# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v1

@@ 5 Experimental Verifications
Experiment setup. We now verify our analysis in practical settings. We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023). We first finetune a GPT-2-large-774M model (Radford et al., 2019) on $5$k samples of PKU-SafeRLHF-QA, and obtain the SFT model. We adopt the GPT2-large-harmless model (Yang et al., 2024) as the ground-truth reward oracle. All experiments are repeated for $3$ seeds. Please see Appendix D for more details.
Implementation details. For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset. For approximate optimization, we instead compute the empirical BT loss. We adopt a pairwise variant of policy gradient instead of PPO to improve training stability: $\mathcal{L}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1}),\hat{r}_{\theta}(y_{2}))\right]^{2}$. During deployment, the ground-truth reward will be scaled by a coefficient $r_{\textup{margin}}$. Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
@@
Verifications of Section 4. We train offline DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al. (2024), and we train on two types of preference: “better” and “safer”, and down-sampled the corresponding training datasets to $1$k-$9$k samples. For DPO training, we directly train the SFT model using DPO; while for pure reward learning, we replace the projection matrix of the SFT model with a linear head. The models are trained under the same setting, and achieve at least $85\%$ training accuracy. Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks.
Figure 2: Experimental Results for Condition 1. We experiment on different reward scales $0.4,1,4$, and observe that the advantage of RL over online DPO is larger for larger reward scales.
Figure 3: Experimental Results for Conditions 2, 3 and 4. The first two figures align with our claims. And the gap in the last figure can be attributed to the mis-specified reward model being too weak.
Figure 4: Experimental Results on Statistical Efficiency. We experiment on two preference types, and pure reward learning is shown to be more data-efficient than surrogate reward learning.
6 Related Works
@@
Nika et al. (2024) provided sub-optimality upper bounds for RLHF and DPO when assuming linear reward class and log-linear policy class, with the un-regularized value as performance metric. Three cases were studied: ➀ realizable ground-truth reward and exact optimization, ➁ realizable ground-truth reward but approximate optimization, as well as ➂ non-realizable reward and exact optimization. Let $n$ be the size of the fixed dataset and $d$ be the feature dimension. For case ➀, both algorithms have a policy bias due to the un-regularized metric, while RLHF has an additional $\Theta(\sqrt{d/n})$ statistical error and that for DPO is $\Theta(d/(\beta n))$. For case ➁, RLHF and DPO both obey a linear convergence to statistical errors and policy biases when using projected gradient descent. For case ➂, aside from statistical errors and policy biases, RLHF has an extra approximation error between the ground-truth reward and best achievable reward, while DPO has an extra bias between the optimal regularized policy and the ideal optimal regularized policy.
Now we focus on comparing with the most relevant prior study, Nika et al. (2024). First, unlike Nika et al. (2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: ➀ it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; ➁ it can help circumvent the unavoidable upper bound of policy bias in the unregularized version. Second, we provide a fine-grained analysis of different model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n\rightarrow+\infty$, and our results are not limited to linear reward and log-linear policy model classes. Third, we improve the statistical analysis of Nika et al. (2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\mathcal{O}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
A more general list of related works is deferred to Appendix A.
@@ 7 Conclusion
This paper provides a fine-grained analysis of the performance gap between two-stage and direct approaches to preference-based policy learning. We theoretically demonstrate a dichotomy of RLHF and DPO under different mis-specification scenarios, and further reveal a implicit representation gap induced by statistical efficiency. Our claims are supported by empirical experiments on LMs.
It is also important to acknowledge our limitations. ➀ While we identify a limitation of training reward models based on BT model, we do not provide a theoretically grounded and practically effective alternative. ➁ Due to computational constraints, our experiments are limited to small-scale models. We hope our insights can motivate the community to further investigate these directions.
@@ Acknowledgement
SSD acknowledges the support of NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P. Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship. RZ and MF acknowledge the support of NSF TRIPODS II DMS-2023166. The work of MF was supported in part by awards NSF CCF 2212261 and NSF CCF 2312775. RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
@@
[…]
Appendix A Supplementary Related Works
Reinforcement learning from human feedback (RLHF). Seminal contributions that showcased RLHF’s applicability to LLMs include foundational work by Christiano et al. [2017], and subsequent research focusing on tasks such as summarization [Stiennon et al., 2020], instruction following [Ouyang et al., 2022], question answering using web-retrieved information [Nakano et al., 2021], and broader AI alignment objectives [Bai et al., 2022]. Theoretical studies of RLHF include pessimism in policy learning [Zhu et al., 2023], overoptimization [Zhu et al., 2024, Liu et al., 2024c], online RLHF [Xiong et al., 2024, Song et al., 2024], robustness [Mandal et al., 2025], and reward models [Wang et al., 2024, Razin et al., 2025b, Huang et al., 2025, Yao et al., 2025].
Direct preference optimization (DPO). There is a rich literature studying offline [Rafailov et al., 2024, Feng et al., 2024], iterative [Dong et al., 2024, Liu et al., 2024a], and online [Guo et al., 2024, Tajwar et al., 2024, Ding et al., 2024, Shi et al., 2025, Chen et al., 2025, Feng et al., 2025] DPO. There are other DPO-style algorithms to directly optimize the policy model from preference signals, such as $\Psi$-PO [Azar et al., 2023], RSO [Liu et al., 2024b], RS-DPO [Khaki et al., 2024], CPO [Xu et al., 2024a], SimPO [Meng et al., 2024], XPO [Xie et al., 2024], VPO [Cen et al., 2024], and OAIF [Guo et al., 2024].
Appendix B Bonus: How Can We Better Model Reward From Preference Signals?
@@
Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section C.12):
@@
Comparing Equation 3 with Equation 5, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 5 to further align with Equation 3. And this will induce an additional coefficient $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ on the data distribution $\mu(y,y^{\prime})$. And this by-product explains that PILAF sampler (a variant of online sampler, see Definition 1) is introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\propto\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$. If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it is a process reward model. However, it is worth noting that the model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2. Therefore, when faced with representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
To alleviate this issue, the remedy could be to learn the preference with alternative modeling approaches to circumvent the BT model setting, which has already be revealed in the successes of Sun et al. [2025], Calandriello et al. [2024]. For example, we can look into the training objective of online IPO [Calandriello et al., 2024, Zhou et al., 2025] (see detailed calculations in Section C.12):
@@
where $\rho(\theta)$ is an online sampling distribution, and it thus can optimize a $\ell_{2}$ distance in an online way. The classification model deployed in Sun et al. [2025] is also promising. We leave this interesting direction for future exploration, and now come back to two-stage RLHF and DPO.
Appendix C Omitted Proofs
C.1 Proof of Proposition 1
@@ C.2 Proof of Theorem 2
$\nabla_{\theta}\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ $=-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]$
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{2\mathsf{sg}\left(Z(\theta)\right)}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left(1+\epsilon_{y,y^{\prime}}\right)\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;,$
@@
$|\epsilon_{y,y^{\prime}}|$ $=\left|\frac{\sigma^{\prime\prime}(\xi_{y,y^{\prime}})}{\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))}\right|\cdot|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|$
$\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime}(R_{\max}+\delta)}\;,$
@@
$Z(\theta):=\sum_{y,y^{\prime}\in\mathcal{Y}}\pi_{\theta}(y)\pi_{\theta}(y^{\prime})/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;.$
@@
$\nabla_{\theta}V_{r^{\star}}^{\pi_{\theta}}$ $=\nabla_{\theta}\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(y)-\hat{r}_{\theta}(y)]$
@@
$V_{r^{\star}}^{\pi_{\theta}}$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}-\frac{1}{4\beta}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;.$
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\theta)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z(\theta)\right)}\left\{-V_{r^{\star}}^{\pi_{\theta}}+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\right\}\;.$
C.3 Proof of Proposition 3

### v2

@@ 5 Experimental Verifications
Experiment setup. We now verify our analysis in practical settings. We consider one common dataset, PKU-SafeRLHF (Ji et al., 2023). We first fine-tune a GPT-2-large-774M model (Radford et al., 2019) on 55k samples of PKU-SafeRLHF-QA, and obtain the SFT model. We adopt the GPT2-large-harmless model (Yang et al., 2024) as the ground-truth reward oracle. All experiments are repeated for $3$ seeds. Please see Appendix C for more details.
Implementation details. For exact optimization, we compute the exact BT loss using the ground-truth reward oracle for each pair in the DPO training dataset. For approximate optimization, we instead compute the empirical BT loss. We adopt a pairwise regression surrogate instead of PPO to improve training stability: $\mathcal{L}_{\textup{RL}}(\theta)=\mathbb{E}_{y_{1},y_{2}\sim\mathsf{sg}\left(\pi_{\theta}\right)}\left[(r(y_{1})-r(y_{2}))-(\hat{r}_{\theta}(y_{1})-\hat{r}_{\theta}(y_{2}))\right]^{2}$. During deployment, the reward score will be scaled by a coefficient $r_{\textup{margin}}$. Besides, since PILAF sampler (see Definition 1) is very close to purely online sampler when $\beta=0.1$, we directly sample $y_{1},y_{2}\sim\pi_{\theta}$ in the implementation of online DPO.
@@
Verifications of Section 4. We train DPO and reward learning on PKU-SafeRLHF, following the practice of Zhou et al. (2024). We train on two types of preference: “better” and “safer”, and down-sample the corresponding training datasets to 11k-99k samples. For DPO training, we directly train the SFT model using DPO; while for pure reward learning, we replace the projection matrix of the SFT model with a linear head. The models are trained under the same setting, and all achieve at least $85\%$ training accuracy. Results are shown in Figure 4, demonstrating that as the number of samples decreases, reward learning outperforms surrogate reward learning across two tasks.
Figure 2: Experimental Results for Condition 1. Experiments with different reward scales $\{0.4,1,4\}$ align with Theorem 2: as the reward scale increases, the second-order deviation in the online DPO objective grows, giving RLHF a clear advantage.
Figure 3: Experimental Results for Conditions 2, 3 and 4. The first two plots (Conditions 2 and 3) are consistent with Propositions 3 and 5. The gap in the last plot can be attributed to the mis-specified reward model being too weak.
Figure 4: Experimental Results on Statistical Efficiency. We experiment on two preference types. Pure reward learning is shown to be more data-efficient than surrogate reward learning.
6 Related Work
Reinforcement learning from human feedback (RLHF). Seminal contributions that showcased RLHF’s applicability to LLMs include foundational work by Christiano et al. (2017), and subsequent research focusing on tasks such as summarization (Stiennon et al., 2020), instruction following (Ouyang et al., 2022), question answering using web-retrieved information (Nakano et al., 2021), and broader AI alignment objectives (Bai et al., 2022). Theoretical studies of RLHF include pessimism in policy learning (Zhu et al., 2023), overoptimization (Zhu et al., 2024; Liu et al., 2024c), online RLHF (Xiong et al., 2024; Song et al., 2024), robustness (Mandal et al., 2025), and reward models (Wang et al., 2024; Razin et al., 2025b; Huang et al., 2025; Yao et al., 2025).
Direct preference optimization (DPO). There is a rich literature studying offline (Rafailov et al., 2024; Feng et al., 2024), iterative (Dong et al., 2024; Liu et al., 2024a), and online (Guo et al., 2024; Tajwar et al., 2024; Ding et al., 2024; Shi et al., 2025; Chen et al., 2025; Feng et al., 2025) DPO. There are other DPO-style algorithms to directly optimize the policy model from preference signals, such as $\Psi$-PO (Azar et al., 2023), RSO (Liu et al., 2024b), RS-DPO (Khaki et al., 2024), CPO (Xu et al., 2024a), SimPO (Meng et al., 2024), XPO (Xie et al., 2024), VPO (Cen et al., 2024), and OAIF (Guo et al., 2024).
@@ 6 Related Work
Nika et al. (2024) provided sub-optimality upper bounds for RLHF and DPO when assuming linear reward class and log-linear policy class, with the un-regularized value as performance metric. Three cases were studied: 1) realizable ground-truth reward and exact optimization, 2) realizable ground-truth reward but approximate optimization, as well as 3) non-realizable reward and exact optimization. Let $n$ be the size of the fixed dataset and $d$ be the feature dimension. For case 1, both algorithms have a policy bias due to the un-regularized metric, while RLHF has an additional $\Theta(\sqrt{d/n})$ statistical error and that for DPO is $\Theta(d/(\beta n))$. For case 2, RLHF and DPO both obey a linear convergence to statistical errors and policy biases when using projected gradient descent. For case 3, aside from statistical errors and policy biases, RLHF has an extra approximation error between the ground-truth reward and best achievable reward, while DPO has an extra bias between the optimal regularized policy and the ideal optimal regularized policy.
Here, we focus on comparing with the most relevant prior study, Nika et al. (2024). First, unlike Nika et al. (2024) which chooses the un-regularized value function as the performance metric, we adopt the regularized version for two reasons: 1) it is the shared original optimization goal of RLHF and DPO, so our choice is to ensure fairness; 2) it can help circumvent the unavoidable upper bound of policy bias in the unregularized version. Second, we provide a fine-grained analysis of different model mis-specifications under exact optimization, i.e., more detailed comparative analysis on reward approximation error and $\mathcal{O}(\mathsf{KL}\left(\pi_{\theta_{\textup{DPO}}}\|\pi^{\star}\right))$ when $n\rightarrow+\infty$, and our results are not limited to linear reward and log-linear policy model classes. Third, we improve the statistical analysis of Nika et al. (2024) on DPO ($\Theta(d_{P}/n)$) and RLHF ($\Theta(\sqrt{d_{R}/n})$), and show that even when $d_{P}=d_{R}=d$ and under realizability assumption, there can still be a large gap between DPO ($\Omega(d/n)$) and RLHF ($\tilde{\mathcal{O}}(\sqrt{k\log d/n})$) where $k\ll d$ is the parameter sparsity.
@@ 7 Conclusion
This paper provides a fine-grained analysis of the performance gap between two-stage and direct approaches to preference-based policy learning. We theoretically demonstrate a dichotomy of RLHF and DPO under different mis-specification scenarios, and further reveal an implicit representation gap induced by statistical efficiency. Our claims are supported by empirical experiments on LMs.
It is also important to acknowledge our limitations. 1) While we identify a limitation of training reward models based on BT model, we do not provide a theoretically grounded and practically effective alternative. 2) Due to computational constraints, our experiments are limited to small-scale models. We hope our insights can motivate the community to further investigate these directions.
@@ Acknowledgement
SSD acknowledges the support of NSF IIS 2110170, NSF DMS 2134106, NSF CCF 2212261, NSF IIS 2143493, NSF IIS 2229881, Alfred P. Sloan Research Fellowship, and Schmidt Sciences AI 2050 Fellowship. RZ acknowledges the support of NSF TRIPODS II DMS-2023166. RS thanks Yunzhen Yao and Yue Wu for insightful discussions.
@@
Appendix A Bonus: How Can We Better Model Reward From Preference Signals?
@@
Following the policy gradient theorem [Sutton et al., 1999], the gradient of this new objective is (see detailed calculations in Section B.11):
@@
Comparing Equation 4 with Equation 6, a natural idea is to apply Taylor’s expansion to extract the $\sigma(\cdot)$ in Equation 6 to further align with Equation 4. And this will induce an additional coefficient $\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ on the data distribution $\mu(y,y^{\prime})$. And this by-product explains that PILAF sampler (a variant of online sampler, see Definition 1) is introduced to align the distorted distribution $\tilde{\mu}(y,y^{\prime})\;\propto\;\mu(y,y^{\prime})\cdot\sigma^{\prime}(r_{\phi}(y)-r_{\phi}(y^{\prime}))$ with $\pi_{\theta^{\star}(r_{\phi})}$. If the reward model is a surrogate reward model, then we can directly deploy PILAF sampler or online sampler; while if it is a pure reward model, then we can implement PILAF sampler or online sampler through logit mixing [Shi et al., 2024, Xu et al., 2025] only when it can provide token-level reward information. However, it is worth noting that model mis-specification can lead the second-order Taylor remainder to be extremely large, as shown in Theorem 2. Therefore, when faced with a representation gap, it could be beneficial to train the (surrogate) reward model on a distribution close to PILAF sampler but is still limited.
To alleviate this issue, we could learn the preference with alternative modeling approaches to circumvent the BT model setting, which has already shown success in Sun et al. [2025], Calandriello et al. [2024]. For example, we can look into the training objective of online IPO [Calandriello et al., 2024, Zhou et al., 2025] (see detailed calculations in Section B.11):
@@
where $\rho(\theta)$ is an online sampling distribution, and it thus can optimize an $\ell_{2}$ distance in an online way. The classification model deployed in Sun et al. [2025] is also promising. We leave this interesting direction for future exploration.
Appendix B Omitted Proofs
Note that in this section, we omit all prompts without loss of generality. For the constructive proof, we can set the number of states to $1$; for the other proofs, we can simply sum over different prompts to extend them to the general case.
B.1 Proof of Proposition 1
@@ B.2 Proof of Theorem 2
$\quad\nabla_{\theta}\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta})$
$=-\beta\underset{y,y^{\prime}\sim\pi^{\mathsf{s}}}{\mathbb{E}}\left[\nabla_{\theta}\log\pi_{\theta}(y)-\nabla_{\theta}\log\pi_{\theta}(y^{\prime})\right]\cdot\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\cdot\left[(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\right]$
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta})$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{1}{2\mathsf{sg}\left(Z_{\theta}\right)}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left(1+\epsilon_{y,y^{\prime}}\right)\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;,$
@@
$|\epsilon_{y,y^{\prime}}|$ $=\left|\frac{\sigma^{\prime\prime}(\xi_{y,y^{\prime}})}{\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))}\right|\cdot|(r^{\star}(y)-r^{\star}(y^{\prime}))-(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))|\leq\frac{\delta}{6\sqrt{3}\sigma^{\prime}(R_{\max}+\delta)}\;,$
@@
$Z_{\theta}:=\sum_{y,y^{\prime}\in\mathcal{Y}}\pi_{\theta}(y)\pi_{\theta}(y^{\prime})/\sigma^{\prime}(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime}))\;.$
@@
$\quad\nabla_{\theta}\left[\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\right]$ (7)
$=\nabla_{\theta}\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(y)-\hat{r}_{\theta}(y)]$
@@
$\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)$ ${\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}-\frac{1}{4\beta}\underset{y,y^{\prime}\sim\pi_{\theta}}{\mathbb{E}}\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\;.$
@@
$\mathcal{L}_{\textup{DPO}}^{\textup{online}}(\pi_{\theta}){\stackrel{{\scriptstyle\scriptstyle\nabla}}{{=}}}\frac{2\beta}{\mathsf{sg}\left(Z_{\theta}\right)}$ $\Bigg\{-\left[\underset{y\sim\pi_{\theta}}{\mathbb{E}}[r^{\star}(x,y)]-\beta\mathsf{KL}\left(\pi_{\theta}\|\pi_{{\mathsf{ref}}}\right)\right]$
$+\frac{1}{4\beta}\underset{y,y^{\prime}\sim\mathsf{sg}\left(\pi_{\theta}\right)}{\mathbb{E}}\left[\epsilon_{y,y^{\prime}}\cdot\left[\left(r^{\star}(y)-r^{\star}(y^{\prime})\right)-\left(\hat{r}_{\theta}(y)-\hat{r}_{\theta}(y^{\prime})\right)\right]^{2}\right]\Bigg\}\;.$
B.3 Proof of Proposition 3